OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 2x − 3y = 3 5x − 4y = 4 −1 0 x can be any number as there are infinitely many solutions to this system There is no x value as there is no solution to this system

OpenStudy (anonymous):

2x − 3y = 3 5x − 4y = 4 2x - 3y - 2x = 3 -3y = 3 - 2x y = (2/3)x - 1 5x - 4y = 4 5x - 4((2/3)x - 1) = 4 5x - (8/3)x + 4 = 4 (15/3)x - (8/3)x + 4 = 4 (7/3)x + 4 = 4 (7/3)x = 0 x=0